OpenStudy (anonymous):

BOYLE'S LAW - MATH EQUATIONS pV = k p1V1 = p2v2 Gas = PiVi = PfVf Initial Pressure = PfVf / Vi Initial Volume = PfVf / Pi Pressure = PiVi / Vf Volume = PiVi / Pf If a gas has a volume of 1 L at a pressure of 270 kPa, what volume will it have when the pressure is increased to 540 kPa? In this case, the equation would be: 270(V) = 540. To double the pressure, I must use long division? 270(1)/540 270(540)/1 ? 540(270)/1 ? 540(1)/270 ? Until somebody actually browses the Chemistry category, I absolutely support my previous indignation that this question belongs in the ma

OpenStudy (anonymous):

I am assuming your first equation pV=k, is pressure times volume equals a constant. In which case, plug in the p and V you are initially given to find that constant. Then for a new pressure, you can use the constant to get the new corresponding volume.

OpenStudy (anonymous):

If you are using the second equation, it is 270*1 = 540*v2. So solving for v2, v2 = 270/540. Long division is perfectly good, although I prefer to reduce to prime factorization.. That is, 270 = 9*3*10 and 540 = 9*3*2*10 So 270/540 is?

OpenStudy (anonymous):

\[270 \div 540 = \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

Which is also 50%. In terms of decimals, 0.50

OpenStudy (anonymous):

right. cool